### v2

@@ Appendix B Experiment Details
In Tables 1 and 2, the names , , and are used for , , and respectively as we only consider the RPS game. The LLAMA-2 70B Chat model is also quantized for runtime efficiency. In Tables 3 and 9 the names , , and are used for , , and respectively to promote consistency with the action spaces also used in IBS and IPD.
@@
B.1 Example Prompts
In this section, we provide example prompts for each prompting type. In all cases, we tailor our provided example to the Iterated Battle of Sexes (IBS) with actions and in which the agent happens to be playing round 5 of 100. There is a slight change in terminology for Rock, Paper, Scissors (RPS) where it is said that the agents ”receive a score of +1/0/-1” rather than that they ”win 10/8/7/5/0 points”. This is just to accommodate for a negative scale of rewards in this game as we did not feel having the reward as always positive would properly reflect the incentive structure of the game to the LLMs.
![Refer to caption](extracted/6181734/images/lm_prompting.png)
![Refer to caption](extracted/6181734/images/lm_prompting_tom.png)
![Refer to caption](extracted/6181734/images/lm_prompting_tom2.png)
![Refer to caption](extracted/6181734/images/qa_prompting.png)
![Refer to caption](extracted/6181734/images/qa_prompting_tom.png)
![Refer to caption](extracted/6181734/images/cot_prompting.png)
![Refer to caption](extracted/6181734/images/cot_prompting_tom.png)
![Refer to caption](extracted/6181734/images/pi_prompting.png)
![Refer to caption](extracted/6181734/images/pi_prompting_tom.png)
![Refer to caption](extracted/6181734/images/reflection_prompting.png)
![Refer to caption](extracted/6181734/images/reflexion_prompting.png)
![Refer to caption](extracted/6181734/images/reflexion_prompting_tom.png)
![Refer to caption](extracted/6181734/images/cot_prompting_3shot.png)
![Refer to caption](extracted/6181734/images/oracle_prompting.png)
![Refer to caption](extracted/6181734/images/oracle_prompting_max.png)
![Refer to caption](extracted/6181734/images/s2a_prompting.png)
![Refer to caption](extracted/6181734/images/s2a_prompting_tom.png)
![Refer to caption](extracted/6181734/images/social_prompting.png)
B.2 Additional Details on Experiments with Single Action Partners
@@
B.3 Additional Experiments with Simple Adaptive Partners
@@
B.4 Additional Experiments on Action Space Inductive Bias
![Refer to caption](extracted/6181734/images/mistral_ipd_tit4tat_tiny.png)
The Good and Bad of Inductive Bias. So far we have used the neutral action representations , , and to avoid contamination following prior work (Binz & Schulz, 2023; Akata et al., 2023). However, it is interesting given our results so far to get a better sense of the degree that prior knowledge in the LLMs impacts performance and the choice of action representation is great way for us to control the extent that this knowledge is evoked. For the best performing models LLAMA-3 70B Instruct and Mistral Large 2, we conduct experiments for all three games playing with both single action and tit for tat style partners. We provide comprehensive results for functional theory of mind performance in Figures 20, 22, 24, 26, 28, 30, 32, 34, 36, 38, 40, and 42, and for literal theory of mind performance in Figures 21, 23, 25, 27, 29, 31, 33, 35, 37, 39, 41, and 43. Here we compare neutral actions with the actual canonical action names for the game, the neutral actions with 20 repetitions to take up a longer portion of the context, and nonsense word actions (see Appendix B). We generally find that both LLM models experience systematic bias that prevents them from converging to optimal performance as the number of interactions grow. Indeed, only LLAMA-3 playing IBS with tit for tat partners (Figure 28) seems on the road to slow convergence. Meanwhile, literal theory of mind seems to be converging in a greater number of settings (Figures 29, 33, 35, and 37). We find that real actions often help with functional theory of mind performance early in the interaction stream while becoming harmful for converged performance as interaction goes on (Figures 19, 20, 26, 30, 32, and 36).
![Refer to caption](extracted/6181734/images/llama3_rps_regret.png)
![Refer to caption](extracted/6181734/images/llama3_rps_acc.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_regret.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_acc.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_regret.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_acc.png)
![Refer to caption](extracted/6181734/images/llama3_rps_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_rps_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_rps_regret.png)
![Refer to caption](extracted/6181734/images/mistral_rps_acc.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_regret.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_acc.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_regret.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_acc.png)
![Refer to caption](extracted/6181734/images/mistral_rps_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_rps_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_acc_tit4tat.png)

### v3

@@ Appendix C Experiment Details
In Tables 1 and 2, the names , , and are used for , , and respectively as we only consider the RPS game. The LLAMA-2 70B Chat model is also quantized for runtime efficiency. In Tables 3 and 10 the names , , and are used for , , and respectively to promote consistency with the action spaces also used in IBS and IPD.
@@
C.1 Example Prompts
In this section, we provide example prompts for each prompting type. In all cases, we tailor our provided example to the Iterated Battle of Sexes (IBS) with actions and in which the agent happens to be playing round 5 of 100. There is a slight change in terminology for Rock, Paper, Scissors (RPS) where it is said that the agents “receive a score of +1/0/-1” rather than that they “win 10/8/7/5/0 points”. This is just to accommodate for a negative scale of rewards in this game as we did not feel having the reward as always positive would properly reflect the incentive structure of the game to the LLMs.
C.2 Additional Details on Experiments with Single Action Partners
@@
C.3 Additional Experiments with Simple Adaptive Partners
@@
C.4 Additional Experiments on Action Space Inductive Bias
The Good and Bad of Inductive Bias. So far we have used the neutral action representations , , and to avoid contamination following prior work (Binz & Schulz, 2023; Akata et al., 2023). However, it is interesting given our results so far to get a better sense of the degree that prior knowledge in the LLMs impacts performance and the choice of action representation is great way for us to control the extent that this knowledge is evoked. For the best performing models LLAMA-3 70B Instruct and Mistral Large 2, we conduct experiments for all three games playing with both single action and tit for tat style partners. We provide comprehensive results for functional theory of mind performance in Figures 20, 22, 24, 26, 28, 30, 32, 34, 36, 38, 40, and 42, and for literal theory of mind performance in Figures 21, 23, 25, 27, 29, 31, 33, 35, 37, 39, 41, and 43. Here we compare neutral actions with the actual canonical action names for the game, the neutral actions with 20 repetitions to take up a longer portion of the context, and nonsense word actions (see Appendix C). We generally find that both LLM models experience systematic bias that prevents them from converging to optimal performance as the number of interactions grow. Indeed, only LLAMA-3 playing IBS with tit for tat partners (Figure 28) seems on the road to slow convergence. Meanwhile, literal theory of mind seems to be converging in a greater number of settings (Figures 29, 33, 35, and 37). We find that real actions often help with functional theory of mind performance early in the interaction stream while becoming harmful for converged performance as interaction goes on (Figures 19, 20, 26, 30, 32, and 36).